1. A 50-kg barbell is lifted at a height of 1.4m from the floor in 1.2 seconds. What is the work done? how much power is expended?


2. A 60 kg block at rest was pushed by a 100 N force on a frictionless horizontal.

A. what is the kinetic energy of the block at a distance of 8 meters?

B. what is the velocity of the blockat a distance of 8 meters?


1
Expert's answer
2022-02-22T04:41:13-0500

Given:

"m=50\:\rm kg"

"h=1.4\:\rm m"

"t=1.2\;\rm s"


The work done

"W=mgh=\rm 50*9.8*1.4=686\: J"

The amount of power

"P=\frac{W}{t}=\frac{686\:\rm J}{1.2\:\rm s}=572\: \rm W"

Given:

"m=60\:\rm kg"

"F=100\:\rm N"

"d=8\;\rm m"


The change in kinetic energy is equal to work done

"\frac{mv^2}{2}=W=Fd=\rm 100*8=800\: J"

The velocity

"v=\sqrt{\frac{2W}{m}}=\sqrt{\frac{2*800}{60}}=5.2\: \rm m/s"